Circular No. 8433

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 YM_{159} (LINEAR-CATALINA) The Catalina Sky Survey has reported observations of two short-tailed

comet suspects on four CCD frames obtained over a 39-min span on Nov. 10.5 UT (observer R. Hill; 0.68-m Schmidt telescope). The head of the slightly fainter of the two was situated about 102′′ west and 33′′ north of that of the brighter and close to end of the latter’s tail. Assuming that the two orbits differed only in T, B. G. Marsden, Harvard-Smithsonian Center for Astrophysics, found that the objects — if real — had to be intermediate-period comets some 4.3–4.4 AU from the earth. The assumption also revealed likely LINEAR observations of a single asteroidal object on Oct. 8 and 24, and a three-night linkage then showed identity with the LINEAR asteroidal object 2003 YM_{159}, observed on 2003 Dec. 17 and 30 (see MPS 109905) — the identity clearly being with the brighter 2004 Nov. 10 object, now designated component A. Discovery observations:

2003/4 UT  α_{2000}  δ_{2000}  Mag. Observer
A  Dec. 17.26424 5^{h}19^{m}48.^{s}87 +37^{o}08′16′′.5 19.1 LINEAR
A  Nov. 10.46635 9 33 16.44 +34 59 27.7 17.6 Hill
B 10.46635 9 33 08.06 +35 00 02.4 18.1 _{”}

The 2004 astrometry, the following orbital elements for component A, and an ephemeris are given on MPEC 2004-V52. T for component B will occur ∼ 0.23 day later.

Epoch = 2005 Mar. 11.0 TT

T = 2005 Feb. 28.5311 TT  ω = 87.^{o}6590
e = 0.444960  Ω = 47.8605 }2000.0
q = 4.411276 AU  i = 19.3584

a = 7.947668 AU n^{o} = 0.0439890 P = 22.41 years

SUPERNOVA 2004fb IN ESO 340-G7 N. Morrell, M. Hamuy, and G. Folatelli, Las Campanas Observatory

and Carnegie Supernova Project, report that a CCD spectrogram (range 360–900 nm), obtained on Oct. 24.09 UT with the Magellan II (Clay) telescope (+ LDSS-2 spectrograph), reveals that SN 2004fb (IAUC 8420) is a type-II supernova showing a blue continuum with strong absorption lines and a prominent P-Cyg profile in Hα. The velocity obtained for the absorption minimum of Hβ is 6620 km/s, assuming the NED host-galaxy recession velocity of 6100 km/s.

2004 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green